If the consumers income rises from rs 10000 to 12000. As a result demand for goods increases from 20 units to 30 units per week. Find the income elasticity of demand per goods

1
Expert's answer
2021-12-15T11:36:25-0500

The income elasticity of demand is calculated as

"E_y=\dfrac{Q_2-Q_1}{Y_2-Y_1}\times \dfrac{Y_1+Y_2}{Q_1+Q_2}\\[0.3cm]"

Therefore

"E_y=\dfrac{30-20}{12000-10000}\times \dfrac{12000+10000}{30+20}\\[0.3cm]
E_y=\boxed{2.2}"